# First Comparative Exoplanetology Within a Transiting Multi-planet System: Comparing the atmospheres of V1298 Tau b and c

### v1

@@
11email: s.barat@uva.nl 22institutetext: School of Earth and Planetary Sciences (SEPS), National Institute of Science Education and Research (NISER), Jatani, India 33institutetext: Astrophysics Reseach Center (ARCO), The Open University of Israel, Ra’anana, 43107, Israel 44institutetext: Frontier Research Institute for Interdisciplinary Sciences, Tohoku University, 6-3 Aramaki aza Aoba, Aoba-ku, Sendai, Miyagi 980-8578, Japan 55institutetext: Department of Geophysics, Graduate School of Science, Tohoku University, 6-3 Aramaki aza Aoba, Aoba-ku, Sendai, Miyagi 980-8578, Japan 66institutetext: Institute of Space and Astronautical Science, Japan Aerospace Exploration Agency, 3-1-1 Yoshinodai, Chuo-ku, Sagamihara, Kanagawa 252-5210, Japan 77institutetext: Cluster for Pioneering Research, RIKEN, 2-1 Hirosawa, Wako, Saitama 351-0198, Japan 88institutetext: Department of Astronomy & Astrophysics, University of California, Santa Cruz, CA 95064, USA. 99institutetext: Department of Astronomy and Astrophysics, University of Chicago, Chicago, IL, USA 1010institutetext: School of Earth and Space Exploration, Arizona State University, Tempe, AZ 85287, USA 1111institutetext: Department of Physics, University of Warwick, Coventry CV4 7AL, UK 1212institutetext: Centre for Exoplanets and Habitability, University of Warwick, Coventry CV4 7AL, UK. 1313institutetext: Lowell Observatory, 1400 W Mars Hill Road, Flagstaff, AZ, 86001, USA 1414institutetext: Center for Computational Astrophysics, Flatiron Institute, New York, NY 10010, USA. 1515institutetext: Astrobiology Center, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan. 1616institutetext: National Astronomical Observatory of Japan,2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan. 1717institutetext: Department of Astronomy, Graduate University for Advanced Studies (SOKENDAI), Tokyo, 2-21-1 Osawa, Mitaka, Japan.
@@
The V1298 Tau system (20-30Myr), is a benchmark young multi-planet system that provides the opportunity to perform comparative exoplanetology between planets orbiting the same star right after their formation. We present the first atmospheric comparison between two planets in the same transiting system: V1298 Tau b and V1298 Tau c. We derive constraints on the mass of planet b and c (¡20M⊕ at 3 confidence level and respectively) and atmospheric metallicity (logZ/Z⊙=-2.04, -0.16 respectively) from atmospheric retrievals. The V1298 Tau planets, are likely to be similar in terms of mass at the current age, implying that both planets are potential sub-Neptune/super-Earth progenitors. However, planet c is expected to lose a higher fraction of its mass compared to planet b given its close proximity to the host star. Alternatively, the observed spectrum of planet c can be explained by atmospheric hazes, which is in contrast to planet b where efficient haze formation can be ruled out. Higher haze formation efficiency in planet c could be due to differences in atmospheric composition, temperature and higher UV flux incident compared to planet b.
@@ 1 Introduction
Demographic studies on mature exoplanets (Fulton et al., 2017; Van Eylen et al., 2018; Szabó & Kiss, 2011) suggest that early evolutionary processes like photoevaporation (Owen & Wu, 2017) and core-powered mass loss (Ginzburg et al., 2018; Gupta & Schlichting, 2019) significantly influence the structure and composition of super-Earths and sub-Neptunes. Therefore, observing planets in their early stages is key to understanding these processes.
@@
V1298 Tau, a 20-30 Myr old weak-lined T-Tauri star in the foreground of the Taurus-Auriga star forming region (Oh et al., 2017; Luhman, 2018), is particularly valuable for such studies. It hosts four transiting planets, making it one of the youngest known multi-planet systems. The inner three planets (V1298 Tau c, d, and b) are in a near 3:2:1 mean-motion resonance (Feinstein et al., 2022; Sikora et al., 2023; David et al., 2019b), suggesting possible migration within a gas disk (Terquem & Papaloizou, 2018; Kajtazi et al., 2023). The orbital period of the outermost planet, V1298 Tau e, remains unconfirmed but is under investigation (Damasso et al., 2023).
@@
In this paper, we compare the atmospheric composition and chemistry of two planets in the V1298 Tau system: V1298 Tau b (0.9 RJ, 24.14 day period) and V1298 Tau c (0.5 RJ, 8.24 day period). A comparison between the known physical properties of both planets is shown in Table 1. V1298 Tau c receives four times the stellar irradiation as V1298 Tau b due to their orbital distance ratio of 1:2. Stellar XUV flux is believed to play an important role in atmospheric chemistry (Kawashima & Ikoma, 2018; Tsai et al., 2023; Venot & Agúndez, 2015; Shulyak et al., 2020) as well as early evolution (Owen & Wu, 2013, 2017), which we aim to test with two planets in the same young system.
@@
Mass measurement for these young planets is challenging due to stellar activity affecting radial velocity (RV) observations (Blunt et al., 2023). Initial RV estimates (Suárez Mascareño et al., 2021), which reported Jovian masses for the two outer V1298 Tau planets, has been challenged (Sikora et al., 2023; Finociety et al., 2023). Barat et al. (2024) estimated V1298 Tau b’s mass to be less than 23M⊕ at 3 level of confidence from HST/WFC3 transmission spectra, and ruled out earlier RV measurements at 5 level of confidence. Ongoing Transit timing variations (TTVs) studies (private communication) suggest that both V1298 Tau b and c are in the sub-Neptune/super-Earth mass regime, indicating they are likely sub-Neptune/super-Earth progenitors with similar masses.
@@
We have observed a transmission spectrum of V1298 Tau c using the same instrumental configuration as Barat et al. (2024) to allow for comparison. We discuss the data reduction, light curve analysis and atmospheric modelling in Section 2, followed by the main results in Section 3. We discuss and interpret our findings in Section 4 and 5.
2 Data analysis and modelling
@@ 2.1 Observations
We observed one primary transit of V1298 Tau c (18th October 2021) with 8 HST orbits using the WFC3 instrument for Program GO 16462. The observations were taken with the G141 grism in bidirectional spatial scanning mode covering a range of 1.1-1.7 m. We used the pixel subarray and SPARS25, NSAMP=5 readout mode which resulted in 88.4 s exposures. We used a 0.23”/second scan rate resulting a 170 pixel long image in the spatial direction.
@@ 2.2 Data reduction and light curve analysis
We use a custom data reduction pipeline for analyzing the raw HST spatially scanned images (Arcangeli et al., 2018; Jacobs et al., 2022; Barat et al., 2024) mirroring the approach for planet b in Barat et al. (2024).
The broadband integrated light curve is shown in Figure 4. It shows ‘hook’ like systematics, which are typical for similar HST observations (Wakeford et al., 2016). Our observations covered only one pre-transit orbit and no ingress/egress observations. Typically, the first orbit of a visit for HST is known to exhibit large systematics, and it is common practice to not include it in time series analysis (e.g see Kreidberg et al., 2014; Jacobs et al., 2022; Barat et al., 2024). However, since we have only one pre-transit orbit, we do not remove it in our light curve analysis. Further details of the light curve fitting methods are provided in Appendix B.
2.3 Atmospheric modelling with ATMO and planet mass
We interpret the transmission spectra of V1298 Tau c and V1298 Tau b using a chemically consistent ATMO retrieval framework (Tremblin et al., 2015; Drummond et al., 2016; Goyal et al., 2018). ATMO has been previously used to analyze various exoplanet atmospheres (see for e.g Evans et al., 2018; Goyal et al., 2019; Carter et al., 2020). Details about the ATMO models are provided in Appendix D.
Figure 1 shows the best-fit atmospheric models for both planets alongside the observed spectra. The retrieved parameters are listed in Table 2, with posterior distributions in Figures 12, 13. Following Barat et al. (2024), we include planet mass as a free parameter. For planet b, the posterior distribution peaks around 8M⊕ with a 3 upper limit of 20M, consistent with Barat et al. (2024). For planet c, the posterior peaks around 17M⊕ with a 3 upper limit of 50M⊕, consistent with previous RV upper limits (Finociety et al., 2023; Sikora et al., 2023). These results suggest both planets are likely in the sub-Neptune/super-Earth mass range.
Our retrievals indicate low atmospheric metallicity for both planets, with planet b having 0.1-0.01 solar and planet c consistent with solar. The temperature of planet b peaks around 850-900K (See Appendix E for further discussion on planet b’s temperature), while planet c’s temperature is unconstrained. No constraints were found on the C/O ratio or clouds for either planet.
We also performed retrievals with fixed masses to assess their impact on other parameters. For V1298 Tau c, we used 10 and 50M, and for planet b, we used 10 and 23M. The posterior distributions for these fixed mass retrievals are shown in Figures 15, 14, with retrieved parameters in Table 2. The overall conclusion of sub-solar/solar atmospheric metallicity for both planets remains valid even with fixed masses.

### v2

@@
11email: s.barat@uva.nl 22institutetext: School of Earth and Planetary Sciences (SEPS), National Institute of Science Education and Research (NISER), Jatani, India 33institutetext: Astrophysics Research Center (ARCO), The Open University of Israel, Ra’anana, 43107, Israel 44institutetext: Frontier Research Institute for Interdisciplinary Sciences, Tohoku University, 6-3 Aramaki aza Aoba, Aoba-ku, Sendai, Miyagi 980-8578, Japan 55institutetext: Department of Geophysics, Graduate School of Science, Tohoku University, 6-3 Aramaki aza Aoba, Aoba-ku, Sendai, Miyagi 980-8578, Japan 66institutetext: Institute of Space and Astronautical Science, Japan Aerospace Exploration Agency, 3-1-1 Yoshinodai, Chuo-ku, Sagamihara, Kanagawa 252-5210, Japan 77institutetext: Cluster for Pioneering Research, RIKEN, 2-1 Hirosawa, Wako, Saitama 351-0198, Japan 88institutetext: Department of Astronomy & Astrophysics, University of California, Santa Cruz, CA 95064, USA. 99institutetext: Department of Astronomy and Astrophysics, University of Chicago, Chicago, IL, USA 1010institutetext: School of Earth and Space Exploration, Arizona State University, Tempe, AZ 85287, USA 1111institutetext: Department of Physics, University of Warwick, Coventry CV4 7AL, UK 1212institutetext: Centre for Exoplanets and Habitability, University of Warwick, Coventry CV4 7AL, UK. 1313institutetext: Lowell Observatory, 1400 W Mars Hill Road, Flagstaff, AZ, 86001, USA 1414institutetext: Center for Computational Astrophysics, Flatiron Institute, New York, NY 10010, USA. 1515institutetext: Astrobiology Center, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan. 1616institutetext: National Astronomical Observatory of Japan, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan. 1717institutetext: Department of Astronomical Science, The Graduate University for Advanced Studies, SOKENDAI, 2-21-1 Osawa, Mitaka, Tokyo 181-8588, Japan.
@@
The V1298 Tau system is a multi-planet system that provides the opportunity to perform comparative exoplanetology between planets orbiting the same star. Because of its young age (20-30 Myr), this system also provides the opportunity to compare the planet’s early evolutionary properties, right after their formation. We present the first atmospheric comparison between two transiting exoplanets within the same multiple planet system: V1298 Tau b and V1298 Tau c. We observed one primary transit for each planet with the Hubble Space Telescope (HST), using Grism 141 (G141) of Wide Field Camera 3 (WFC3). We fit the spectroscopic light curves using state-of-the-art techniques to derive the transmission spectrum for planet c and adopt the transmission spectrum of planet b obtained with the same observing configuration and data analysis methods from previous studies. We measure the mass of planet b and c (, M⊕; respectively) from the transmission spectrum and find the two planets to have masses in the Neptune/sub-Neptune regime. Using atmospheric retrievals, we measure and compare the atmospheric metallicities of planet b and c (logZ/Z⊙=-2.04, logZ/Z⊙= -0.16, respectively), and find them consistent with solar/sub-solar, which is low (at least 1 order or magnitude) compared to known mature Neptune/sub-Neptune planets. This discrepancy could be explained by ongoing early evolutionary mechanisms, which are expected to enrich the atmospheres of such young planets as they mature. Alternatively, the observed spectrum of planet c can be explained by atmospheric hazes, which is in contrast to planet b, where efficient haze formation can be ruled out. Higher haze formation efficiency in planet c could be due to differences in atmospheric composition, temperature and higher UV flux compared to planet b. In addition, planet c is likely to experience a higher fraction of mass loss compared to planet b, given its proximity to the host star.
@@ 1 Introduction
Demographic studies on mature exoplanets (Szabó & Kiss 2011; Fulton et al. 2017; Fulton & Petigura 2018; Van Eylen et al. 2018) suggest that early evolutionary processes like photoevaporation (Owen & Wu 2017) and core-powered mass loss (Ginzburg et al. 2018; Gupta & Schlichting 2019) significantly influence the structure and composition of low mass planets, such as Neptunes and sub-Neptunes. Therefore, observing planets right after their formation is the key to understanding these early evolutionary processes.
V1298 Tau, a 20-30 Myr old weak-lined T-Tauri star in the foreground of the Taurus-Auriga star-forming region (Oh et al. 2017; Luhman 2018), is particularly valuable for such studies. It has a temperature of 5000 K, with a radius and mass of 1.340.05 R⊙ and 1.10.05 M⊙, respectively, and has a spectral type between K0-K1 (David et al. 2019a). It is a pre-main sequence star, and shows rotational variability at the level of 2% peak-to-peak and a rotational period of 2.860.01 days (David et al. 2019a; Feinstein et al. 2022; Sikora et al. 2023). Such rotational variability is typical of young pre-main sequence stars (e.g. see David et al. 2016; Mann et al. 2016, 2022) and is due to the presence of spots and faculae on their photospheres. Feinstein et al. (2021) reproduced the optical photometric variability using a model with 20% spot coverage on the surface assuming a temperature contrast of 500 K between the photosphere and the spots.
@@
V1298 Tau hosts four transiting planets, making it one of the youngest known transiting multi-planet systems. The inner three planets (V1298 Tau c, d, and b) are in a near 3:2:1 mean-motion resonance (David et al. 2019b; Feinstein et al. 2022; Sikora et al. 2023), suggesting possible migration within a gas disk (Terquem & Papaloizou 2018; Kajtazi et al. 2023). The mutual inclinations between the planets in this system are known to be low (Johnson et al. 2022), indicating planar orbital geometry for this system. The orbital period of the outermost planet, V1298 Tau e, remains unconfirmed but is under investigation (Damasso et al. 2023).
With a wealth of observations of exoplanet atmospheres, we have entered an era of understanding atmospheric processes, planet formation and early evolution through comparative studies. There have been previous studies aimed at understanding the physical and chemical processes of mature transiting exoplanet atmospheres from a statistical point of view. For example, the dependence of thermal structure and opacity sources in the atmosphere on insolation (e.g, see Garhart et al. 2020; Baxter et al. 2020), emergence of clouds in the atmosphere (Sing et al. 2016; Fu et al. 2017; Keating et al. 2019), disequilibrium chemistry (Baxter et al. 2021; Fortney et al. 2020; Tsai et al. 2023; Baeyens et al. 2022) and the formation and composition of hazes (Gao et al. 2020; Brande et al. 2024) has been studied by comparing the atmospheres of different exoplanets. But, differences in the host stars can introduce biases due to uncertainty in their elemental abundances, activity levels and evolution history. These challenges can be mitigated by performing comparative exoplanetology between multiple planets in the same system, since they share their host star. In this paper, we compare the atmospheric metallicity, chemistry and potential evolution of two planets in the V1298 Tau system: V1298 Tau b (0.9 RJ, 24.14 day period) and V1298 Tau c (0.5 RJ, 8.24 day period) by studying their atmospheres. Given its young age, this system also allows for atmospheric comparison with mature transiting exoplanets as well as, widely separated directly imaged planets which have similar age. A comparison between the known physical properties of planet b and c is shown in Table 1.
@@
V1298 Tau c receives four times the stellar irradiation as V1298 Tau b due to their orbital distance ratio of 1:2. Stellar XUV flux is believed to play an important role in atmospheric chemistry (Venot & Agúndez 2015; Kawashima & Ikoma 2018; Shulyak et al. 2020; Tsai et al. 2023) as well as early evolution (Owen & Wu 2013, 2017), which we aim to test with two planets in the same young system.
@@
Mass measurement for these young planets is challenging due to stellar activity induced jitter affecting radial velocity (RV) observations (Brems et al. 2019; Blunt et al. 2023). Initial RV studies (Suárez Mascareño et al. 2021) reported Jovian masses (20070, 400100 M⊕) for the two outer V1298 Tau planets (b and e, respectively) and mass upper limits (70 and 100 M⊕) for the inner two (c and d, respectively). Follow-up RV observations of this system have challenged these mass measurements (Sikora et al. 2023; Finociety et al. 2023). Finociety et al. (2023) reports an upper limit of 50 M⊕ for planet c and d and an upper limit of 150 M⊕ for planet b. Sikora et al. (2023) reports a mass of 2010 M⊕ for planet c and upper limits of 36 and 160 M⊕ for planets d and b, respectively. The follow-up RV papers are consistent with Suárez Mascareño et al. (2021) for the mass of V1298 Tau e, but do not agree with the mass of planet b from Suárez Mascareño et al. (2021). Barat et al. (2024) estimated V1298 Tau b’s mass to be less than 23 M⊕ at 3 level of confidence from HST/WFC3 transmission spectra, and ruled out RV estimate from Suárez Mascareño et al. (2021) at 5 level of confidence. Ongoing Transit timing variations (TTVs) studies (Livingston et al in prep) suggest that both V1298 Tau b and c are in the sub-Neptune mass regime.
@@
In this paper, we compare the transmission spectra of V1298 Tau b and c. For planet b, we adopt the spectrum obtained in Barat et al. (2024). We analyze the primary transit observations of V1298 Tau c, observed using the same instrumental configuration as Barat et al. (2024), and present the transmission spectrum of planet c. We discuss the data reduction and light curve analysis for the visit of V1298 Tau c in Section 2, followed by the main results in Section 3. We discuss and interpret our findings in Section 4 and 5.
| Parameter | V1298 Tau b | V1298 Tau c |
| Period [days] 11111111Sikora et al. (2023) | 24.1404100.000022 | 8.2487200.000024 |
| Semi-major axis [AU] | 0.17160.0028 | 0.08390.0014 |
| Equilibrium temperature [K] | 68515 | 97921 |
| Radius [R⊕] | 9.950.37 | 5.240.24 |
| Known mass [M⊕] | 23 22222222Barat et al. (2024) | 2010 11111111Sikora et al. (2023) |
2 Data analysis and modeling
@@ 2.1 Observations
We observed one primary transit of V1298 Tau c (18th October 2021) with 8 HST orbits using the WFC3 instrument for Program GO 16462. The observations were taken with the G141 in bidirectional spatial scanning mode covering a range of 1.1-1.7 m. We used the pixel subarray and SPARS25, NSAMP=5 readout mode, which resulted in 88.4 s exposures. We used a 0.23”/second scan rate, resulting in a 170 pixel long image in the spatial direction. The observations for V1298 Tau b had been taken using the same instrumental settings, with 10 HST orbits. See Barat et al. (2024) for further details.
@@ 2.2 Data reduction and light curve analysis
We use a custom data reduction pipeline for analyzing the raw HST spatially scanned images (Arcangeli et al. 2018, 2019; Jacobs et al. 2022; Barat et al. 2024) for the V1298 Tau c visit. The same pipeline had been used for analyzing the visit of planet b in Barat et al. (2024). We outline the data reduction methods we use in Appendix A.
The broadband integrated white light curve for the visit of V1298 Tau c is shown in Figure 1. It shows ‘hook’ like systematics for each orbit, which have been observed for similar HST observations (Berta et al. 2012; Deming et al. 2013; Ranjan et al. 2014). We also see that the orbital ramp is much larger in the first orbit compared to the rest of the orbits for this visit. This is a typical feature of HST visits where the first orbit shows larger systematics (Wakeford et al. 2016; Arcangeli et al. 2021). The light curves also show a long-term curvature, which is likely due to the rotational variability from active regions on the photosphere of this young star. This variability has been reported by Barat et al. (2024) and is well known for young pre-main sequence stars which have large spot coverage and fast rotation rates (e.g. see David et al. 2016, 2019b; Mann et al. 2022). Such variability is known to affect multi-epoch transmission spectra (e.g, Désert et al. 2011b; Berta et al. 2012; Rackham et al. 2019).
These observations of planet c were planned using linear ephemeris from Feinstein et al. (2022), but this system is known to exhibit TTVs and the actual transit occurred 2 hours prior to the predicted time from the linear ephemeris. Therefore, our observations covered only one pre-transit orbit and no exposures during ingress or egress. Typically, the first orbit of a visit for HST is known to exhibit large systematics, and it is common practice not to include it in time series analysis (e.g. see Berta et al. 2012; Ranjan et al. 2014; Jacobs et al. 2022; Barat et al. 2024). In the case of V1298 Tau b, there were three pre-transit orbits, and therefore, the first orbit could be removed from the analysis. However, since we have only one pre-transit orbit for the visit of V1298 Tau c, we could not remove the first orbit while analyzing its light curve.
The white light curve of planet c has been modeled assuming a physically motivated charge-trapping model RECTE (Zhou et al. 2017) to model the ‘hook-like’ orbital ramps and a second order polynomial function to model the baseline for the planet c visit. The RECTE model is known to fit the first orbit better compared to other known methods (e.g, Berta et al. 2012). The planet transit was modeled using batman (Kreidberg 2015), with orbital parameters taken from Sikora et al. (2023). This same approach had been used to fit the white light curve of planet b (Barat et al. 2024).
A linear limb darkening model was assumed for V1298 Tau. Barat et al. (2024) fit for the linear limb darkening coefficient for the visit of planet b. Since the planet c visit does not have any ingress/egress coverage, we have used the fitted limb darkening for V1298 Tau from Barat et al. (2024) as a fixed parameter for the analysis of the planet c visit (Figure 10). V1298 Tau b and c have impact parameters (0.34 and 0.46, respectively) consistent within 1 of each other (David et al. 2019b, a). Furthermore, Johnson et al. (2022) report a relatively low mutual inclination (019∘), indicating a planar orbital geometry of the V1298 Tau system. Therefore, both planets would transit a similar part of the stellar photosphere, and are expected to to show similar limb darkening during their primary transits. Further details of the light curve fitting methods are provided in Appendix B.
We obtain 17 spectroscopic light curves for the planet c visit by using 7-pixel bins. The same binning had been used for planet b in Barat et al. (2024). For the analysis of spectroscopic light curves of planet c, the semi-major axis is fixed from the white light curve fit of planet c. This approach had also been followed for the planet b visit. In the analysis of planet b (Barat et al. 2024), the common-mode method (using white light curve systematics models to correct the spectroscopic light curves) was applied to de-trend the spectroscopic light curves, assuming that the systematics were achromatic. However, the first orbit is known to exhibit wavelength-dependent systematics (Zhou et al. 2017, 2020), and therefore in the analysis of V1298 Tau c we could not use the common-mode technique for de-trending spectroscopic light curves. We used RECTE (Zhou et al. 2017) to model the ‘hook-like’ orbital ramps and a second order polynomial function to model the baseline of the spectroscopic light curves for the planet c visit. The parameters for the RECTE as well as the polynomial coefficients, were fitted for each spectroscopic light curve. We tested different models for the baseline (linear, quadratic and cubic). We found that the quadratic model fits the observed baseline the best. See Appendix C for further details on the baseline model. The limb darkening coefficients were fixed from Barat et al. (2024).
We derive the transmission spectrum of V1298 Tau c following the approach outlined in this Section, and we present it in Figure 2, along with the spectra of planet b, which has been adopted from Barat et al. (2024).
2.3 Atmospheric models from ATMO
We interpret the transmission spectrum of V1298 Tau c using the equilibrium chemistry ATMO retrieval framework (Tremblin et al. 2015; Drummond et al. 2016; Goyal et al. 2018). In this paper, for a consistent comparison with planet c, we also reanalyze the spectrum of V1298 Tau b adopted from Barat et al. (2024), using the same framework as applied for planet c. ATMO has been previously used to analyze various exoplanet atmospheres (see for e.g Evans et al. 2018; Goyal et al. 2019; Carter et al. 2020). ATMO’s python based retrieval framework was applied and benchmarked in Lewis et al. (2020) and Rathcke et al. (2021) for HAT-P-41b and WASP-79b, respectively. This same retrieval framework is applied to V1298 Tau b and c in this work. This framework includes a Nested sampling Bayesian sampler DYNESTY (Speagle 2020) coupled to ATMO.
We model the planets using a 1D atmosphere with an isothermal T-P profile and calculate the molecular abundances, assuming chemical equilibrium. Atmospheric metallicity and C/O ratio are used as parameters to calculate the elemental abundances. For a given metallicity, in our model the C/O ratio is implemented by keeping the number of C atoms fixed, and adjusting the number of O atoms. We include a grey opacity source to model clouds. We include H2-H2, H2-He Collision Induced opacity (CIA), as well as H2O, CH4 and CO high temperature molecular opacity in the retrieval. These molecules are chosen, as the HST bandpass includes absorption features from these molecules. We include planet mass as a free parameter in our retrieval setup for both planets b and c. We put similar uniform priors on atmospheric metallicity (0.01 - 1000 solar), O/H ratio (0.01 - 100 solar), degree of cloudiness (0.1 - 100 Rayleigh scattering cross-section at 350 nm) for both V1298 Tau b and c. We put uniform prior on the mass between 5 - 50 M⊕ for planet c (Finociety et al. 2023; Sikora et al. 2023) and between 5-25 M⊕ for planet b (Barat et al. 2024).